OpenStudy (anonymous):

HELP! True? If as x approaches x0 of lim(f g)(x)=0 then f(x)=0 if approaches x-->x0 or g(x)=o as it approaches x---> x0?

OpenStudy (helder_edwin):

is your question something like this: if \[ \large \lim_{x\to x_0}(f\circ g)(x)=0 \] then \[ \large \lim_{x\to x_0}f(x)=0\quad\text{or}\quad\lim_{x\to x_0}g(x)=0 \]

OpenStudy (anonymous):

yes exactly!

OpenStudy (anonymous):

is this true?

OpenStudy (helder_edwin):

give me a second

OpenStudy (helder_edwin):

it is false

OpenStudy (helder_edwin):

here it is the complete justification: let \(f(x)=x^2-2\) and \(g(x)=x+1\) then \[ \large (f\circ g)(x)=f(g(x))=f(x+1)=(x+1)^2-2 \] then \[ \large (f\circ g)(x)=x^2+2x-1 \]

OpenStudy (helder_edwin):

ok so far?

OpenStudy (helder_edwin):

r u there @Compgroupmail ?

OpenStudy (anonymous):

I am here. I'm back now.

OpenStudy (helder_edwin):

can u tell why the proposition if false from what i already did?

OpenStudy (anonymous):

yes, I understand now. THanks :) !!

OpenStudy (anonymous):

I'll add you as a fan/

OpenStudy (helder_edwin):

well anyway. if u solve \((f\circ g)(x)=0\) u get \[ \large x=-1\pm\sqrt{2} \] then \[ \large \lim_{x\to-1+\sqrt{2}}(f\circ g)(x)=0 \] but \[ \large \lim_{x\to-1+\sqrt{2}}f(x)\neq0\quad\text{and} \lim_{x\to-1+\sqrt{2}}g(x)\neq0 \]